OpenStudy (anonymous):

The motion of a particle is given by a=6v^1/2, where a is in feet per second^2 and v in feet per second. when t=0, s=6 ft and v=0. Find the relations between v and t, s and t, v and s

OpenStudy (anonymous):

you know that a=dv/dt \[\frac{ dv }{ \sqrt{v} } =6dt \] integrate this then you can get the relation between v and t .. then you can calculate other relations using v=ds/dt

OpenStudy (anonymous):

sir the \[\sqrt{v} = 0\] ?? or not equal to zero?

OpenStudy (anonymous):

sorry sir im not really good at integrating .. specially with squareroot

OpenStudy (anonymous):

no it is not.. you should check your math ..

OpenStudy (anonymous):

oh how did you come up with \[\frac{ dv }{ \sqrt{v} }=6dt\]

OpenStudy (anonymous):

you should put dv/dt in a..

OpenStudy (anonymous):

thanks sir :)